# The Geometry of Codes for Random Access in DNA Storage

### v1

@@ 1 Introduction
This paper is organized as follows. Section 2 provides the definitions used throughout the paper, formally introduces the Random Access Problem, and offers a brief overview of existing results on the topic. In Section 3, we introduce a geometric object in the projective plane, which we call a balanced quasi-arc. Intuitively, a balanced quasi-arc is a set of projective points in which three specific points, referred to as fundamental points, are in the span of many subsets of other points in the set, more so than any of the non-fundamental points. This characteristic of the set of points increases the likelihood of recovering these fundamental points when sampling uniformly at random from the set, ultimately leading to codes that perform well regarding the Random Access Problem. To the best of our knowledge, the codes we obtain based on this construction outperform any of the existing constructions of codes in terms of the Random Access Problem. All of this will be discussed in detail in Section 4. In Section 5, we investigate the parameters of a class of codes with a rate of $1/2$, first constructed in [2], that have a random access expectation strictly smaller than their dimension $k$. By closely examining these parameters, we are able to resolve a conjecture proposed in [2] that states the ratio of their random access expectation to $k$ is strictly smaller than $0.9456k$ as $k$ tends to infinity. The approach we take is inspired by a result from [1] and it gives a new way of showing that the random access expectation is smaller than $k$. Finally, in Section 6 we conclude the paper and propose some possible directions for future research.
@@ 2 Preliminaries
    $[n]$ denotes the set of integers from $1$ to $n$, i.e., $[n]:=\{1,\ldots,n\}$;
@@
Whenever the $k$ information strands are encoded using a generator matrix $G\in\mathbb{F}_{q}^{k\times n}$ one can think of an encoded strand as its corresponding column in the matrix $G$ and recovering the $i$-th information strand corresponds to recovering the $i$-th standard basis vector, that is, it should belong to the span of the already recovered columns of $G$. This justifies that when we care about recovering the $i$-th information strand, we only care about the set of columns of the matrix $G$, and not the order in which they appear. Moreover, the set of columns can be seen as a set of vectors in $\mathbb{F}_{q}^{k}$ and hence points in $\mathrm{PG}(k-1,q)$, the $(k-1)$-dimensional projective space over $\mathbb{F}_{q}$, as we care only about their span. We recall the definition of $\mathrm{PG}(k-1,q)$ and of other related objects that we use throughout this paper.
@@
The projective geometry $\mathrm{PG}(k-1,q)$ is the projective space of dimension $k-1$ over a finite field $\mathbb{F}_{q}$ with $q$ elements. It is defined as the set of equivalence classes of the non-zero vectors in $\mathbb{F}_{q}^{k}$, where two vectors are considered equivalent if they differ by a scalar multiple.
@@
We start by surveying the results from [2] contributing to the random access problem. Among other things, it was shown that by connecting to the coupon collector’s problem [22, 23, 24, 25], if the $k$ information strands are encoded by an MDS code, then the expected number of reads to decode all $k$ information strands is $H_{n}-H_{n-k}$, where $H_{i}$ is the $i$-th harmonic number. This result is optimal for minimizing the expected number of reads. However, analyzing MDS codes for the random access problem, it was proven that the expected number of reads is $k$, which is also what an identity code would give. Moreover, in [2] it was shown that codes with expectation $ck$, for $c<1$, exist. In Section 5, we examine in more detail the construction of a class of codes with rate $1/2$ that has a low random access expectation, thereby proving an open conjecture proposed by the authors of [2].
As a step towards a better understanding of the properties that give a good code in terms of having low random access expectation, in [1] a general formula for the expectation was given. Let $\mathscr{G}=\{P_{1},\dots,P_{n}\}\subseteq\mathrm{PG}(k-1,q)$ be a multiset of points of rank $k$. For $P\in\mathscr{G}$ and $11\leq s\leq n-$ 1, let
@@ 3 Balanced Quasi-Arcs in Projective Planes
Let $\mathscr{G}\subseteq\mathrm{PG}(2,q)$ be a set of points, let :=$\mathscr{F}:=\{E_{1},E_{2},E_{3}\}$ be a set of non-collinear points in $\mathrm{PG}(2,q)$ and let $x\geq 0$ be an integer. We call $\mathscr{G}$ a balanced quasi-arc of weight $x$ if there exists a partition of $\mathscr{G}$ of the form $\mathscr{G}=\mathscr{F}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$ such that
@@
Roughly speaking, a balanced quasi-arc of weight $x$ behaves like an arc of the projective plane with respect to all the lines apart from the fundamental lines, which are $(x+2)$-secant lines (i.e. lines that intersects $\mathscr{G}$ in $x+2$ points). A balanced quasi-arc of weight 00 corresponds to a set of three non-collinear points in $\mathrm{PG}(2,q)$.
@@
Observe that a balanced quasi-arc of weight $x$ in $\mathrm{PG}(2,q)$ has size $33x+$ 3 and the notion of fundamental points is well-defined as $E_{1},E_{2}$ and $E_{3}$ are the only points of $\mathscr{G}$ lying on two $(x+2)$-secant lines.
@@
The $(S_{1},S_{2},S_{3})$-projective triangle from Definition 3.4 is inspired by the projective triangle which, in our terminology, is a $(S,S,N)$-projective triangle in PG$(2,q)$ where $q$ is odd, $S$ is the set of squares in $\mathbb{F}_{q}$, and $N$ is the set of non-squares in $\mathbb{F}_{q}$. The projective triangle turns out to be a minimal blocking set in $\mathrm{PG}(2,q)$; A minimal blocking set in $\mathrm{PG}(2,q)$ is a point set, where every line of $\mathrm{PG}(2,q)$ meets this set in at least one point and the set does not contain any line, and which is minimal with respect to this property and set-theoretic inclusion. When $q$ is an odd prime, Blokhuis in [26] proved that the size of a minimal blocking set is at least $33(q-1)/2+$ 3 and so the projective triangle attains the lower bound with equality; see also [27].
@@
Let $H$ be a subgroup of $(\mathbb{F}_{q}^{*},\cdot)$ and suppose there exists $\tilde{H}\subset\mathbb{F}_{q}^{*}\setminus H$ such that $|H|=|\tilde{H}|=x$. Then the $(H,H,\tilde{H})$-projective triangle is a balanced quasi-arc of weight $x$.
@@
Note that the points of the $(H,H,\tilde{H})$-projective triangle $\mathscr{G}$ can be written as
@@
where :=$\mathscr{F}:=\{E_{1},E_{2},E_{3}\}$, $\mathscr{G}_{1}=\{(0:1:-s_{1})\colon s_{1}\in H\}$, $\mathscr{G}_{2}=\{(-s_{2}:0:1)\colon s_{2}\in H\}$ and $\mathscr{G}_{3}=\{(1:-s_{3}:0)\colon s_{3}\in\tilde{H}\}$. Clearly, this is a partition of $\mathscr{G}$. In order to show that $\mathscr{G}$ is a balanced quasi-arc we need to prove that three points $P_{1}=(0:1:-s_{1})\in\mathscr{G}_{1}$, $P_{2}=(-s_{2}:0:1)\in\mathscr{G}_{2}$ and $P_{3}=(1:-s_{3}:0)\in\mathscr{G}_{3}$, with $s_{1},s_{2},s_{3}\neq 0$, are not collinear. Indeed, $P_{1},P_{2},P_{3}$ are collinear if and only if
@@
Let $q>2$ be a prime power and consider
  • (i)
    $H$ as the subgroup of squares in $(\mathbb{F}_{q}^{*},\cdot)$ and $\tilde{H}=\mathbb{F}_{q}^{*}\setminus H$ if $q$ is odd;
  • (ii)
    $H=(\mathbb{F}_{q/2}^{*},\cdot)$ and $\tilde{H}\subseteq(\mathbb{F}_{q}^{*},\cdot)\setminus H$ with $|\tilde{H}|=(q-1)/2$ if $q$ is even.
We have that the $(H,H,\tilde{H})$-projective triangle is a balanced quasi-arc of weight $\lfloor\frac{q-1}{2}\rfloor$.
@@
    Let $q=5$. Note that $1$ and $4$ are the only squares in $(\mathbb{F}_{5}^{*},\cdot)$, and so we let $H:=\{1,4\}$ and $H^{\prime}:=\{2,3\}$. Then by Corollary 3.7 the $(H,H,\tilde{H})$-projective triangle is a balanced quasi-arc of weight $(q-1)/2=2$. More precisely, $\mathscr{G}_{2}=\mathscr{F}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$ where $\mathscr{F}$ is the set of fundamental points, and
@@
    Let $q=4$ and consider the finite field $\mathbb{F}_{4}$ made of the elements in $\{0,1,\alpha,\alpha+1\}$ where $\alpha$ is a root of the irreducible polynomial $x^{2}+x+1$ over $\mathbb{F}_{q}$. Then we let $H:=\{1\}$, $\tilde{H}:=\{\alpha\}$ and we obtain that $\mathscr{G}_{1}=\mathscr{F}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$ where $\mathscr{F}$ is the set of fundamental points, and
@@
For any $q$ prime power and $x\leq\lfloor\frac{q-1}{2}\rfloor$ a positive integer we have that there exists a balanced quasi-arc in $\mathrm{PG}(2,q)$ of weight $x$.
@@
By Corollary 3.7, we know that there exists a balanced quasi-arc $\mathscr{G}$ in $\mathrm{PG}(2,q)$ of weight $\lfloor\frac{q-1}{2}\rfloor$. By definition, $\mathscr{G}$ can be partitioned into four sets
@@ 4 A Construction for $k=3$
In this section, we show that if we consider the set of points $\mathscr{G}$ as a balanced quasi-arc, we can compute $\alpha_{F}(\mathscr{G},s)$ for all $11\leq s\leq n-$ 1. Besides obtaining (relatively) clean formulas for the values of $\alpha_{F}(\mathscr{G},s)$, we demonstrate that the random access expectation for these sets of points is strictly less than $k$. We separately address the case where the fundamental points appear only once in our sets, and the case where multiplicities are added to the fundamental points. We will also need the following notation.
@@ 4.1 Distinct points
For $23\leq s\leq x+$ 2, the only $s$-sets that do not recover $E_{1}$ are those that only consist of points in $E_{2}E_{3}$. This gives that there are $\binom{3x+3}{s}-\binom{x+2}{s}$ sets that recover $E_{1}$.
@@
$\mathbb{E}[\tau_{F}(\mathscr{G}_{x})]=3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}.$
@@
$=\sum_{s=1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}-\sum_{s=3}^{x+2}\frac{\binom{3x+3}{s}}{\binom{3x+2}{s}}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}-\sum_{s=x+3}^{3x+2}\frac{\binom{3x+3}{s}}{\binom{3x+2}{s}}$
$=\sum_{s=1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}-\sum_{s=3}^{3x+2}\frac{\binom{3x+3}{s}}{\binom{3x+2}{s}}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=\sum_{s=1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}-\sum_{s=3}^{3x+2}\frac{3x+3}{3x+3-s}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=\sum_{s=1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}-\sum_{s=1}^{3x}\frac{3x+3}{s}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=\sum_{s=3x+1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=1+\frac{3x+3}{3x+2}+\frac{3x+3}{3x+1}-\frac{1}{3x+2}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=2+\frac{3x+3}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=2+\frac{3x+3}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}$
$=3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}.$
Note that in the above computations we repeatedly used the fact that for integers $a>b\geq 1$ we have $\binom{a}{b}/\binom{a-1}{b}=a/(a-b)$. ∎
@@
    Let $\mathscr{G}_{2}$ be defined as the balanced quasi-arcs in Example 3.8 (i) of weight 2 over $\mathbb{F}_{5}$. We have $\mathbb{E}[\tau_{F}(\mathscr{G}_{2})]\approx 2.87$.
@@
    Let $\mathscr{G}_{1}$ be defined as the balanced quasi-arcs in Example 3.8 (ii) of weight 1 over $\mathbb{F}_{4}$. We have $\mathbb{E}[\tau_{F}(\mathscr{G}_{2})]=2.9$.
@@
$\mathbb{E}[\tau_{F}(\mathscr{G})]$ $=3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}$
$=3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\sum_{s=3}^{x+2}\prod_{i=2}^{s-1}\frac{x+2-i}{3x+2-i}.$
@@
$\mathbb{E}[\tau_{F}(\mathscr{G}_{x})]$ $\leq 3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\sum_{s=3}^{x+2}\prod_{i=2}^{s-1}\frac{1}{3}$
$=3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\sum_{s=3}^{x+2}\left(\frac{1}{3}\right)^{s}$
$=3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\left(\sum_{s=0}^{x}\left(\frac{1}{3}\right)^{s}-1\right)$
$=3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\left(\frac{1-(1/3)^{x+1}}{1-1/3}-1\right)$
$=3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\left(\frac{1}{2}-\frac{1}{2}\left(\frac{1}{3}\right)^{x}\right).$
@@
$3+\frac{2}{3x+1}-\frac{2(x+2)(x+1)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\left(\frac{1}{2}-\frac{1}{2}\left(\frac{1}{3}\right)^{x}\right)\sim 3-\frac{2}{9}+\frac{1}{18}=3-\frac{1}{6}$

### v2

@@ 1 Introduction
This paper is organized as follows. Section 2 provides the definitions used throughout the paper, formally introduces the Random Access Problem, and offers a brief overview of existing results on the topic. In Section 3, we introduce a geometric object in the projective plane, which we call a balanced quasi-arc. Intuitively, a balanced quasi-arc is a set of projective points in which three specific points, referred to as fundamental points, are in the span of many subsets of other points in the set, more so than any of the non-fundamental points. This characteristic of the set of points increases the likelihood of recovering these fundamental points when sampling uniformly at random from the set, ultimately leading to codes that perform well regarding the Random Access Problem. To the best of our knowledge, the codes we obtain based on this construction outperform any of the existing constructions of codes in terms of the Random Access Problem. All of this will be discussed in detail in Section 4. In Section 5, we investigate the parameters of a class of codes with a rate of $1/2$, first constructed in [2], that have a random access expectation strictly smaller than their dimension $k$. By closely examining these parameters, we are able to resolve a conjecture proposed in [2] that states the ratio of their random access expectation to $k$ is strictly smaller than $0.9456$ as $k$ tends to infinity. The approach we take is inspired by a result from [1] and it gives a new way of showing that the random access expectation is smaller than $k$. Finally, in Section 6 we conclude the paper and propose some possible directions for future research.
@@ 2 Preliminaries
    $[n]$ denotes the set of integers from $1$ to $n$, i.e., $n]:=\{1,\ldots,n\}$;
@@
Whenever the $k$ information strands are encoded using a generator matrix $G\in\mathbb{F}_{q}^{k\times n}$ one can think of an encoded strand as its corresponding column in the matrix $G$ and recovering the $i$-th information strand corresponds to recovering the $i$-th standard basis vector, that is, it should belong to the span of the already recovered columns of $G$. This justifies that when we care about recovering the $i$-th information strand, we only care about the set of columns of the matrix $G$, and not the order in which they appear. Moreover, the set of columns can be seen as a set of vectors in $\mathbb{F}_{q}^{k}$ and hence points in $\mathrm{PG}(k-1,q)$, the (k1)(k-1)-dimensional projective space over $\mathbb{F}_{q}$, as we care only about their span. We recall the definition of $\mathrm{PG}(k-1,q)$ and of other related objects that we use throughout this paper.
@@
The projective geometry $\mathrm{PG}(k-1,q)$ is the projective space of dimension k1k-1 over a finite field $\mathbb{F}_{q}$ with $q$ elements. It is defined as the set of equivalence classes of the non-zero vectors in $\mathbb{F}_{q}^{k}$, where two vectors are considered equivalent if they differ by a scalar multiple.
@@
We start by surveying the results from [2] contributing to the random access problem. Among other things, it was shown that by connecting to the coupon collector’s problem [22, 23, 24, 25], if the $k$ information strands are encoded by an MDS code, then the expected number of reads to decode all $k$ information strands is $n(H_{n}-H_{n-k})$, where $H_{i}$ is the $i$-th harmonic number. This result is optimal for minimizing the expected number of reads. However, analyzing MDS codes for the random access problem, it was proven that the expected number of reads is $k$, which is also what an identity code would give. Moreover, in [2] it was shown that codes with expectation $ck$, for $c<1$, exist. In Section 5, we examine in more detail the construction of a class of codes with rate $1/2$ that has a low random access expectation, thereby proving an open conjecture proposed by the authors of [2].
As a step towards a better understanding of the properties that give a good code in terms of having low random access expectation, in [1] a general formula for the expectation was given. Let $\mathscr{G}=\{P_{1},\dots,P_{n}\}\subseteq\mathrm{PG}(k-1,q)$ be a multiset of points of rank $k$. For $P\in\mathscr{G}$ and $1\leq s\leq n-1$, let
@@ 3 Balanced Quasi-Arcs in Projective Planes
Let $\mathscr{G}\subseteq\mathrm{PG}(2,q)$ be a set of points, let $\mathscr{F}:=\{E_{1},E_{2},E_{3}\}$ be a set of non-collinear points in $\mathrm{PG}(2,q)$ and let $x\geq 0$ be an integer. We call $\mathscr{G}$ a balanced quasi-arc of weight $x$ if there exists a partition of $\mathscr{G}$ of the form $\mathscr{G}=\mathscr{F}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$ such that
@@
Roughly speaking, a balanced quasi-arc of weight $x$ behaves like an arc of the projective plane with respect to all the lines apart from the fundamental lines, which are $(x+2)$-secant lines (i.e. lines that intersects $\mathscr{G}$ in $x+2$ points). A balanced quasi-arc of weight 0 corresponds to a set of three non-collinear points in $\mathrm{PG}(2,q)$.
@@
Observe that a balanced quasi-arc of weight $x$ in $\mathrm{PG}(2,q)$ has size $3x+3$ and the notion of fundamental points is well-defined as $E_{1},E_{2}$ and $E_{3}$ are the only points of $\mathscr{G}$ lying on two $(x+2)$-secant lines.
@@
The $(S_{1},S_{2},S_{3})$-projective triangle from Definition 3.4 is inspired by the projective triangle which, in our terminology, is a $(S,S,N)$-projective triangle in PG(2,q)(2,q) where $q$ is odd, $S$ is the set of squares in $\mathbb{F}_{q}$, and $N$ is the set of non-squares in $\mathbb{F}_{q}$. The projective triangle turns out to be a minimal blocking set in $\mathrm{PG}(2,q)$; A minimal blocking set in $\mathrm{PG}(2,q)$ is a point set, where every line of $\mathrm{PG}(2,q)$ meets this set in at least one point and the set does not contain any line, and which is minimal with respect to this property and set-theoretic inclusion. When $q$ is an odd prime, Blokhuis in [26] proved that the size of a minimal blocking set is at least 3(q1)/2+33(q-1)/2+3 and so the projective triangle attains the lower bound with equality; see also [27].
@@
Let $H$ be a subgroup of $(\mathbb{F}_{q}^{*},\cdot)$ and suppose there exists $\tilde{H}\subset\mathbb{F}_{q}^{*}\setminus H$ such that $|H|=|\tilde{H}|=x$. Then the (H,H,H~)($H,\tilde${H})-projective triangle is a balanced quasi-arc of weight $x$.
@@
Note that the points of the (H,H,H~)($H,\tilde${H})-projective triangle $\mathscr{G}$ can be written as
@@
where $\mathscr{F}:=\{E_{1},E_{2},E_{3}\}$, $\mathscr{G}_{1}=\{(0:1:-s_{1})\colon s_{1}\in H\}$, $\mathscr{G}_{2}=\{(-s_{2}:0:1)\colon s_{2}\in H\}$ and $\mathscr{G}_{3}=\{(1:-s_{3}:0)\colon s_{3}\in\tilde{H}\}$. Indeed, $P_{1},P_{2},P_{3}$ are collinear if and only if
@@
Let $q$ be an odd prime power and consider $H$ as the subgroup of squares in $(\mathbb{F}_{q}^{*},\cdot)$ and $\tilde{H}=\mathbb{F}_{q}^{*}\setminus H$. We have that the (H,H,H~)($H,\tilde${H})-projective triangle is a balanced quasi-arc of weight $\frac{q-1}{2}$.
@@
    Let $q=5$. Note that $1$ and $4$ are the only squares in $(\mathbb{F}_{5}^{*},\cdot)$, and so we let $H:=\{1,4\}$ and $H^{\prime}:=\{2,3\}$. Then by Corollary 3.7 the (H,H,H~)($H,\tilde${H})-projective triangle is a balanced quasi-arc of weight (q1)/2=2(q-1)/2=2. More precisely, $\mathscr{G}=\mathscr{F}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$, where $\mathscr{F}$ is the set of fundamental points, and
@@
    Let $q=4$ and consider the finite field $\mathbb{F}_{4}$ made of the elements in $\{0,1,\alpha,\alpha+1\}$ where $\alpha$ is a root of the irreducible polynomial $x^{2}+x+1$ over $\mathbb{F}_{2}$. Then we let $H:=\{1\}$, $\tilde{H}:=\{\alpha\}$ and we obtain that $\mathscr{G}=\mathscr{F}\cup\mathscr{G}_{1}\cup\mathscr{G}_{2}\cup\mathscr{G}_{3}$, where $\mathscr{F}$ is the set of fundamental points, and
@@
For any $q$ odd prime power and $x\leq\frac{q-1}{2}$ a positive integer we have that there exists a balanced quasi-arc in $\mathrm{PG}(2,q)$ of weight $x$.
@@
By Corollary 3.7, we know that there exists a balanced quasi-arc $\mathscr{G}$ in $\mathrm{PG}(2,q)$ of weight $\frac{q-1}{2}$. By definition, $\mathscr{G}$ can be partitioned into four sets
@@
Remark 3.11.
For the case $q$ even, a possible choice for $H$ is the largest subgroup of the multiplicative group of $(\mathbb{F}_{q}^{*},\cdot)$ and as $\tilde{H}$ a subset of the same size in $\mathbb{F}_{q}^{*}\setminus H$. This clearly depends on the prime factorization of q1q-1; see [28, 29].
@@ 4 A Construction for $k=3$
In this section, we show that if we consider the set of points $\mathscr{G}$ as a balanced quasi-arc, we can compute $\alpha_{F}(\mathscr{G},s)$ for all $1\leq s\leq n-1$. Besides obtaining (relatively) clean formulas for the values of $\alpha_{F}(\mathscr{G},s)$, we demonstrate that the random access expectation for these sets of points is strictly less than $k$. We separately address the case where the fundamental points appear only once in our sets, and the case where multiplicities are added to the fundamental points. We will also need the following notation.
@@ 4.1 Distinct points
For $3\leq s\leq x+2$, the only $s$-sets that do not recover $E_{1}$ are those that only consist of points in $E_{2}E_{3}$. This gives that there are $\binom{3x+3}{s}-\binom{x+2}{s}$ sets that recover $E_{1}$.
@@
$\mathbb{E}[\tau_{F}(\mathscr{G}_{x})]=3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}.$
@@
$=\sum_{s=1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}-\sum_{s=3}^{x+2}\frac{\binom{3x+3}{s}}{\binom{3x+2}{s}}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}-\sum_{s=x+3}^{3x+2}\frac{\binom{3x+3}{s}}{\binom{3x+2}{s}}$
$=\sum_{s=1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}-\sum_{s=3}^{3x+2}\frac{\binom{3x+3}{s}}{\binom{3x+2}{s}}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=\sum_{s=1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}-\sum_{s=3}^{3x+2}\frac{3x+3}{3x+3-s}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
where in the last equality we used the fact that for integers $a>b\geq 1$ we have $\binom{a}{b}/\binom{a-1}{b}=a/(a-b)$. We further simplify to obtain the the above expression is equal to
$\mathbb{E}[\tau_{F}(\mathscr{G}_{x})]$ $=\sum_{s=1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}-\sum_{s=1}^{3x}\frac{3x+3}{s}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=\sum_{s=3x+1}^{3x+3}\frac{3x+3}{s}-\frac{1}{3x+2}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=1+\frac{3x+3}{3x+2}+\frac{3x+3}{3x+1}-\frac{1}{3x+2}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$
$=2+\frac{3x+3}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}$.
Using the definition of the binomial coefficient we obtain that
$\frac{\binom{x+2}{s}}{\binom{3x+2}{s}}=\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}$
and so we further simplify to obtain that
$\mathbb{E}[\tau_{F}(\mathscr{G}_{x})]$ $=2+\frac{3x+3}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}$
$=3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}.\qed$
@@
    Let $\mathscr{G}$ be defined as the balanced quasi-arcs in Example 3.8 (i) of weight 2 over $\mathbb{F}_{5}$. We have $\mathbb{E}[\tau_{F}(\mathscr{G})]\approx 2.87$.
@@
    Let $\mathscr{G}$ be defined as the balanced quasi-arcs in Example 3.8 (ii) of weight 1 over $\mathbb{F}_{4}$. We have $\mathbb{E}[\tau_{F}(\mathscr{G})]=2.9$.
@@
$\mathbb{E}[\tau_{F}(\mathscr{G})]$ $=3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\sum_{s=3}^{x+2}\prod_{i=0}^{s-1}\frac{x+2-i}{3x+2-i}$
$=3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\sum_{s=3}^{x+2}\prod_{i=2}^{s-1}\frac{x+2-i}{3x+2-i}.$
@@
$\mathbb{E}[\tau_{F}(\mathscr{G}_{x})]$ $\leq 3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\sum_{s=3}^{x+2}\prod_{i=2}^{s-1}\frac{1}{3}$
$=3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\sum_{s=3}^{x+2}\left(\frac{1}{3}\right)^{s-2}$
$=3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\left(\sum_{s=0}^{x}\left(\frac{1}{3}\right)^{s}-1\right)$
$=3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\left(\frac{1-(1/3)^{x+1}}{1-1/3}-1\right)$
$=3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\left(\frac{1}{2}-\frac{1}{2}\left(\frac{1}{3}\right)^{x}\right).$
@@
$3+\frac{2}{3x+1}-\frac{2((x+2)(x+1)+x)}{(3x+2)(3x+1)}+\frac{(x+2)(x+1)}{(3x+2)(3x+1)}\left(\frac{1}{2}-\frac{1}{2}\left(\frac{1}{3}\right)^{x}\right)\sim 3-\frac{2}{9}+\frac{1}{18}=3-\frac{1}{6}$